OpenStudy (anonymous):

∫_2^3▒(9x-2x^3 )dx evauluate

OpenStudy (anonymous):

\[∫_2^3▒(9x-2x^3 )dx\]

OpenStudy (akshay_budhkar):

▒ whats this?

OpenStudy (anonymous):

smudge

OpenStudy (anonymous):

if it is nothing then this is straight forward. anti derivative is \[\frac{9}{2}x^2-\frac{1}{2}x^4\]

OpenStudy (anonymous):

\[\int_{2}^{3} \color{red}{\huge{(9x - 2x^3)dx}}\]

OpenStudy (anonymous):

a little latex is a dangerous thing

OpenStudy (anonymous):

Woo latex!

OpenStudy (akshay_budhkar):

why satellite?lol

OpenStudy (anonymous):

then \[\frac{9}{2}x^2-\frac{1}{2}x^4|_2^3\] get -10 i think

OpenStudy (anonymous):

typo it should be x^4

OpenStudy (anonymous):

looks good. yes for sure

OpenStudy (anonymous):

And \(x^{2}\) as well.

OpenStudy (anonymous):

yeah typo

OpenStudy (anonymous):

\[\color{#ff0099}{\huge{\text{the world is a carousel of color}}}\] of color

OpenStudy (akshay_budhkar):

its -10 sorry

OpenStudy (anonymous):

\(\huge{\color{red}{R}\color{orange}{A}\color{yellow}{I}\color{green}{N}\color{blue}{B}\color{indigo}{O}\color{violet}{W}}\)

OpenStudy (anonymous):

\[\color{red}{I}\color{yellow}{S}\color{green}{H}\color{Blue}{AA}N\]

OpenStudy (anonymous):

i go that thanks dall thanks